+sec2

is a conservative vector field (on its implied domain)


a. Find its potential function
b. Find F dr where C is the curve shown below and given by the vector equation

(sin-t-2, cos-t-2, cos(nt) 式t)

Solve using concepts of Vector Fields, Line Integrals, and/or The Fundamental Theorem for Line Integrals.

3 2 2 1 2 1 0 2

+sec2
F dr
(sin-t-2, cos-t-2, cos(nt) 式t)
3 2 2 1 2 1 0 2
Answer #1

af 2x f+sec...3) int partially (1)wrt x er zog(y, z) from (2)and (4) g (y,z) = tan-i (y) +h (z) then f= log(x2+yz)+tan-ı(y)+h